OpenStudy (anonymous):

Which of the following are solutions to the equation sinxcosx=sqrt(3)/4? 1. pi/12 +npi 2. pi/6 +npi 3. pi/12 +npi/2 4. 2pi/6 +npi

hero (hero):

Multiply both sides by 4 to get \(4\sin(x)\cos(x) = \sqrt{3}\) Then square both sides: \((4\sin(x)\cos(x))^2 = (\sqrt{3})^2\)

hero (hero):

\(16\sin^2x\cos^2x = 3\)

hero (hero):

Let \(\cos^2x = 1 - \sin^2x\): \(16\sin^2x(1 - \sin^2x) = 3\) Distribute: \(16\sin^2x - 16\sin^4x = 3\) Rearrange the equation in to the form \(ax^2 + bx + c = 0\) \(-16\sin^4x + 16\sin^2x = 3\) \(-16\sin^4x + 16\sin^2x - 3 = 0\)

hero (hero):

Let \(\sin^2x = y\) \(-16y^2 + 16y - 3 = 0\) \(16y^2 - 16y + 3 = 0\) Solve for y, then substitute \(\sin^2x\) back in to isolate x. Make sure to check your answers.

OpenStudy (anonymous):

Thank you!

hero (hero):

If you run into any problems, let me know.